OpenStudy (anonymous):

Simplify the given expression to rational exponent form, justify each step by identifying the properties of rational exponents used. All work must be shown.

OpenStudy (anonymous):

\[\frac{ 1 }{ 3\sqrt{x^-6} }\]

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

sqrt has exponent of 1/2 exponent in denominator can be written as neg \[\rightarrow \frac{1}{3} (x-6)^{-1/2}\]

OpenStudy (anonymous):

so thats the answer? or is there more to do after that?

OpenStudy (dumbcow):

thats it ...rational exponent just means write exponent as a fraction

OpenStudy (anonymous):

thank you